Night report 2017-09-20

Observers: Zuzanna Kostrzewa-Rutkowska, Simon Hodgkin and Mariusz Gromadzki,  

TAT: Lluís Galbany, Joe Lyman

Support team: Morgan Fraser, Emma Callis

Calibrations: biases, dome arcs+flats, grz twilight flats, 

Conditions:  we start with seeing ~2" then improve to 0.8-1.3" 

Non-detection: AT2017lax

UT       Target            Type            Comment


23:40    AT2017egv        follow-up        Gr11 silt=1.5"

00:20    AT2017lax        class            Non-detection after detailed inspection of aqu image by Simon 

00:25    AT2016cvk        follow-up        Vgri seeing 0.8"

01:10    PG2336           standard photo

01:22    Feige110         standard spec

01:46    ATLAS17kxb       class            Gr13  Ia +9d

02:03    AT2016iks        follow-up        BVgri

02:31    AT2017gut        class            Gr13  blue continuum

02:55    AT2017kwi        class            Gr13  Ia around maximum  

03:14    Feige110         standard spec

03:40    AT2017gah        follow-up        Gr11&Gr16

04:44    PG0231           standard phot

04:58    ATLAS17kwe       class            Gr13  Ia around maximum

05:15    EFOSC2->SOFI

05:30    HIP109059        tell_AT2017gah   BG 

05:45    AT2017gah        follow-up        BG

07:16    AT2017gah        follow-up        JHK

07:50    HIP029337        tell_AT2017gah   BG

07:55    AT2017fzw        follow-up        BG

08:23    AT2017fzw        follow-up        JHK

08:34    EG21             standard         BG

08:47    AT2017gci        follow-up        JHK